Check whether or not the matrix

A=[1 2 3]

[0 2 3]

[0 0 3]

Is diagonalisable. If it is find a matrix P so that P^-1AP is a diagonal matrix. If A is not diagonalisable, obtain its minimal polynomial.





1
Expert's answer
2021-02-03T03:25:31-0500

Check whether or not the matrix


"A=\begin{pmatrix}
 1 & 2 &3\\
 0 &2&3\\0 &0&3
\end{pmatrix}"

Is diagonalisable. If it is find a matrix P so that P^-1AP is a diagonal matrix. If A is not diagonalisable, obtain its minimal polynomial.

find the charateristics polynomial


"f(\lambda) =det(A-\lambda\Iota)=0"

"\begin{vmatrix}
 1-\lambda & 2 &3\\
 0 &2-\lambda&3\\0 &0&3-\lambda
\end{vmatrix} = (1-\lambda)[(2-\lambda)(3-\lambda)]=0"

"\lambda=1 , \lambda-2, \lambda=3"


since we have three eigen values, then the matrix is diagonalisable

when "\lambda=1"

"\begin{pmatrix}
 0 & 2 &3\\
 0 &1&3\\0 &0&2
\end{pmatrix} \begin{pmatrix}
 x_1 \\
 x_2\\x_3 
\end{pmatrix}=0"


"x_2 +3x_3=0......(i)"

"x_2+3x_3=0.......(ii)"

"x_3=0"

"\implies x_1=x_2=x_3=0"


when "\lambda=2"

"\begin{pmatrix}
 -1 & 2 &3\\
 0 &0&3\\0 &0&1
\end{pmatrix} \begin{pmatrix}
 x_1 \\
 x_2\\x_3 
\end{pmatrix}=0"


"-x_1+2x_2+3x_3=0"

"x_3 =0"

since "x_3=0\implies 2x_2=x_1"

since there is no restriction placed on x1

x1=r(arbitrary) where can be any number

let r=2

then x1 = 2

x2 = 1


when "\lambda=3"

"\begin{pmatrix}
 -2 & 2 &3\\
 0 &-1&3\\0 &0&0
\end{pmatrix} \begin{pmatrix}
 x_1 \\
 x_2\\x_3 
\end{pmatrix}=0"


bringing out the augmented matrix


"\begin{pmatrix}
 -2 & 2 &3&0\\
 0 &-1&3&0\\0 &0&0&0
\end{pmatrix}R_1\mapsto R_1-R_2" "\begin{pmatrix}
 -2 &3&0&0\\
 0 &-1&3&0\\0 &0&0&0
\end{pmatrix}"


"\therefore -2x_1+3x_2=0\implies x_1={3 \over 2}x_2"


"-1x_2 +4x_3=0\implies x_2 =4x_3"


since there is no restriction placed on x1

x2=s (arbitrary) where can be any number

let s=2

then x1 = 3


x3 = "{1 \over 2}"


Matrix P ="\begin{pmatrix}
 0 & 2 &3\\
 0 & 1&2 \\0&0&{1 \over 2}
\end{pmatrix}"